Sharon has 30 nickels and dimes having having a total value of $2.35. How many of each kind does she have? I have a chart to fill in, find the equation then solve.
Number of coins x value per coin = total value
Nickels would be x for number of coins times value would be 5 equals total value 5x
Dimes would be 30-X as number of coins times value would be 10 so I"m not sure what total value would be ? Maybe 10(30-x) is that right?
I know we have 30 total coins with a total value of 235. How do I properly set up this equation to solve this. Thank you very much.

Answer by richwmiller(17219)

This belongs in word problems coins
Don't use x for anything here.
Use n for nickels and d for dimes
We need two equations.
What do we know the total number of coins.
What coins do we have nickels and dimes
n+d=30
one equation down.
We need one more
What else do we know ? the total value of the coins. $2.35
But we also know two things that aren't mentioned.
We know the value of each coin.
Each nickel is worth 5 cents and each dime is worth 10 cents
5n+10d=235
Did you even look at any other coin problems? They all work like this.
so now we have two equations
n+d=30
5n+10d=235
Sometimes the problem only asks for one coin's value or the number of only one coin. So be sure to check the question. Here we want to know not the value of the coins but the number of coins.
We know the total is 30
So each one must be less than 30
n+d=30
5n+10d=235
n=30-d
5*(30-d)+10d=235
150-5d+10d=235
150+5d=235
5d=85
d=17
n=13